The manager of a ski resort is working on two special packages. Plan A would give one person 4 nights lodging and 3 dinners for at most $320. Plan B would give one person 5 nights lodging and 4 dinners for at most $406. What is the most one dinner could cost?
A) $62
B) $45.71
C) $45.11
D) $24
I cant figure out how to break this down? THANKS SO MUCH!

Answer by Paul(988)

First equation:
4x+3y=320
Second equation:
5x+4y=406

Solve by elimination method.
Mulitply equation 1 by 4 and equation 2 by 3

4(4x+3y=320)
3(5x+4y=406)


16x+12y=1280 (1)
15x+12y=1218 (2)

Subbtract 1 by 2
x=62

solve for y
4(62)+3y=320
248+3y=320
3y=72
y=24

Hence, about $24--> dinner cost.
Paul.